OpenStudy (chris215):

find the exact value by using a half-angle identity. sin (22.5 degrees)

OpenStudy (asnaseer):

Express \(\cos(2\theta)\) in terms of \(\sin(\theta)\) and then substitute \(\theta=22.5^0\)

OpenStudy (crazypizzalover):

use mathway.com

OpenStudy (chris215):

sqrt (2- sqrt 2)?

OpenStudy (jdoe0001):

\(\bf 22.5\cdot 2\implies 45\qquad thus\qquad \cfrac{45}{2}\iff 22.5 \\ \quad \\ sin\left(\cfrac{{\color{red}{ 45^o}}}{2}\right)=\pm \sqrt{\cfrac{1-cos({\color{red}{ 45^o}})}{2}}\)

OpenStudy (asnaseer):

not quite @chris215 - if you list the steps you took then I can help you spot where you may have made a mistake

OpenStudy (chris215):

using the half angle formula for sin : sin((1/2)x) = 0 +/- sprt (1-cos(x) /2)

OpenStudy (chris215):

then i substitted 45 for x

OpenStudy (chris215):

and got sin(22.5) = sqrt ((1 - sqrt(2)/2))

OpenStudy (asnaseer):

what did you use for the value of cos(45)?

OpenStudy (chris215):

sqrt of 2

OpenStudy (asnaseer):

that is not correct

OpenStudy (asnaseer):

\[\cos(45)=\frac{1}{\sqrt{2}}=\frac{\sqrt{2}}{2}\]

OpenStudy (chris215):

oh yeah thats what I got sorry but next I used the positive sqrt and simplified

OpenStudy (asnaseer):

if you use that value and simplify then you do not end up with what you put above - please list your steps so I can spot where your mistake may have been

OpenStudy (asnaseer):

you can use the draw button below if that makes it easier for you

OpenStudy (chris215):

ok ill try my best to write it out on here

OpenStudy (asnaseer):

e.g.: